OpenStudy (anonymous):

Multiply (sqrt10 +1)(8sqrt10 +1)

OpenStudy (anonymous):

\[(\sqrt10 +1) (8\sqrt10 +1)\]

OpenStudy (anonymous):

multiply out you get \[(8\times 10+\sqrt{10}+8\sqrt{10}+1\] \[81+9\sqrt{10}\]